BG5CONV(1) BG5CONV(1) NAME bg5conv−convertaTeXdocumentinBig5encodinginto‘preprocessed’form. SYNOPSIS bg5conv<infile>outfile DESCRIPTION Big5encodingfortraditionalChineseusesthecharacters{,},and\whichhavespecialmeanings inTEXdocuments. AfterprocessingaLATEX2εdocumentwhichcontainsBig5characterswithbg5convyouneed notcareaboutthesespecialcharacters. ThisfilterispartoftheCJKmacropackageforLATEX2ε. SEEALSO sjisconv(1),cefconv(1),cef5conv(1),cefsconv(1),extconv(1), theCJKdocumentationfiles. AUTHOR WernerLemberg<wl@gnu.org> CJKVersion4.3.0 20-Jun-1999 1